# Retrieving Semantics from the Deep: an RAG Solution for Gesture Synthesis

### v1

@@ 5 Conclusion
In this work, we present RAG-Gesture – a novel approach that leverages interpretable retrieval algorithms to introduce inference-time RAG to a diffusion-based gesture generator. Therefore, it combines advantages of the pure retrieval-based and fully data-driven approaches to synthesize natural yet semantically meaningful gestures. Perceptual evaluation demonstrates that this type of neuro-explicit approach is better than either of the two approaches and existing state-of-the-art methods. Moreover, we demonstrate that LLM-driven gesture type prediction and explicit discourse relations can mitigate the lack of semantic understanding in current gesture generation architectures. Lastly, our framework can be extended to synthesize gestures that exhibit task-specific gestural patterns like referential or emotion-specific gestures.
@@ Acknowledgements.
Table 5: Overall Multi-modality
We first provide additional evaluation of the method in terms of multi-modality and per-speaker analysis. Then we discuss limitations and provide user study details. Lastly, we discuss evaluation metrics and implementation details.
6 Per-speaker FID/Mulitmodality comparison with SOTA
@@
To perform a robust evaluation on speaker generalizability of our framework, we provide per-speaker FID and Multi-modality metrics in Tab. 4. We observe that our framework achieves best FID for large number of the speakers, which shows that our method generalizes well to the speaker specific patterns and idiosyncrasies despite taking no seed gestures at input. CaMN [39], achieves lower FID score overall because it uses seed input from the ground truth data which results in lower scores. Due to the same reason, EMAGE [40] also gets lower FID score.
@@
However, CaMN and EMAGE do not perform well on Multi-modality [12]. This metric computes Euclidean distance between different generated gestures for a single speech input. Since, both of these methods always generate same gestures for a given speech input, they perform worse in terms of Multi-modality, which makes them less ideal for diverse gesture generation. In contrast, our approach gets the best results showing diverse gesture generation capabilities of our model ( Tab. 5).
7 Details on User Study
For the perceptual evaluation of gesture generation capability, 56 participants were shown a randomly sampled set of 16 forced-choice questions. Our study consists of four sections corresponding to four different user studies. First section focuses on the comparison with the state-of-the-art approaches. Each question comprises of a side-by-side animation of our method along with one of EMAGE [40], Audio2Photoreal [51], RemoDiffuse [68] or the ground-truth. Second section is a short section where we compare LLM-driven gesture generation with discourse-based synthesis. Third, we perform pair-wise comparisons between results from one of the RAG baselines (derived from ReMoDiffuse [68]) and our method. Lastly, we ablate different approaches to perform RAG and control its extent.
In first three sections, we try to evaluate naturalness and appropriateness. Specifically, we ask two questions (a) “Which of the two gestures look natural?” and (b) “Which of them looks appropriately aligned to what the person is saying?”. In the last section, we add an additional question which focuses on gauging the semantic appropriateness of gestures in the retrieval window, with a goal of evaluating the RAG capability. We highlight the identified words from the retrieval algorithms and add an additional question: “Which of the two have better gestures in the highlighted section, especially at the capitalized word in the prompt?”
8 Limitations
Our method relies on the sparse semantic data which is extracted from the BEAT2 dataset [40]. This creates a scarcity of good exemplars which can be used during the database matching steps. As a result, the LLM-based Gesture Type sometimes struggles to find good contextual matches for each gesture type and word identified by the LLM.
9 Evaluation Metrics
FID.
Beat Alignment Score.
L1 Divergence.
This metric (also called L1 variance) measures the distance of all frames in a single gesture generation from their mean. It is helpful in identifying synthesized gestures that are static and unexpressive.
Diversity.
It computes the average pairwise Euclidean distance of the gesture generations in the test set.
Multi-modality.
This metric requires sampling different gesture motions for a single speech input from the generative model [12]. Then, it computes Euclidean distance between those generated gestures. It probes the diverse sampling capabilities of a generative model.
10 Implementation Details
@@ 10.1 Input Representations
Representing speech and its transcription is highly important aspect of diffusion-based gesture modelling process. In our experiments, we found that changing the structure of text embeddings affects gesture understanding during the learning process, which consequently is reflected during the synthesis phase as well. To construct our text representation, we build a per-frame embedding with corresponding word embeddings residing on each frame. We compute the word embeddings by combining sub-word token activations from last 4 layers of BERT model [13].
@@ 10.2 Decoupled gesture encoding
We utilize time-aware VAE architecture by Mughal et al. [47]. This architecture utilizes seperate encoders for frame window chunks of original motion to encode each chunk into an encoding. Then, it jointly decodes all of the chunks together to reconstruct the original motion. We use $N=150$ representing $10$ seconds of motion at $15$ frames per second. Moreover, the motion chunk length of our time-aware VAE is $15$, making each chunk encoding correspond to $1$ second of motion. This results in an chunked gesture encoding of length $10$ for each body part. Finally, we concatenate all 4 body part encodings along the time axis with separators in between them, resulting in $M=40+3=43$.
@@ 10.4 Details on LLM Prompting
“You are an expert in human gestures. You need to identify words that may elicit semantically meaningful gestures(deictic, iconic, metaphoric) and their types: (a) Metaphoric Gesture: Represents abstract ideas or concepts physically, creating a vivid mental image. (b) Iconic Gesture: Mimics the shape or action of the object or concept being described. (c) Deictic Gesture: Points to or indicates a person, object, or location. Format your response as a python list of python tuples of (word, type). For example: [(’hello’, ’beat’), (’world’, ’iconic’)]”
@@
Identify at most 2 important words which are more likely to elicit semantically meaningful gestures and what are types of those gestures in following text: “TEXT”.
@@ 10.5 Baseline Retraining Details
To be consistent with single speaker evaluation on BEAT2 dataset, we utilize released model weights by EMAGE [40]. For other approaches (including ours), we retrain the method on single speaker data belong to speaker “scott”. Since there are no available models for the chosen baselines which have been trained on all speakers in BEAT2 dataset, we train all the methods on complete dataset through their provided codebases. Methods which do not contain speaker specific generalizations like Audio2Photoreal [51], are modified to include a speaker embedding along with text and speech embeddings. Moreover, Audio2Photoreal is adapted to support the skeletal format of BEAT2.
Lastly, ReMoDiffuse, originally released for text-to-motion task, is modified for gesture synthesis and their retrieval process is implemented using text feature similarity method. To ablate the training-time RAG, we further modify the ReMoDiffuse architecture to use our gesture encoding and our retrieval algorithms, which is referred to as “ReMoDiffuse+Our Retrieval”. Analysis on this is given in the Ablative Analysis section.

### v2

@@ 5 Conclusion
In this work, we present RAG-Gesture – a novel approach that leverages interpretable retrieval algorithms to introduce inference-time RAG to a diffusion-based gesture generator. Therefore, it combines advantages of the pure retrieval-based and fully data-driven approaches to synthesize natural yet semantically meaningful gestures. Perceptual evaluation demonstrates that this type of neuro-explicit approach is better than either of the two approaches and existing state-of-the-art methods. Moreover, we demonstrate that LLM-driven gesture type prediction and discourse relations can enhance semantic understanding in current gesture generation architectures. Lastly, our framework can be extended to synthesize gestures that exhibit task-specific gestural patterns like referential or emotion-specific gestures.
@@ Acknowledgements.
We first discuss limitations of our method and provide details on the perceptual evaluation. Further, we elaborate on evaluation metrics and provide additional experiments and analyses. Lastly, we discuss our implementation and provide analysis and runtime details for it.
6 Limitations
Our method relies on the sparse semantic data which is extracted from the BEAT2 dataset [40]. This creates a scarcity of good exemplars which can be used during the database matching steps. As a result, the LLM-based Gesture Type algorithm sometimes struggles to find good contextual matches for each gesture type and word identified by the LLM.
7 Details on User Study
For the perceptual evaluation of gesture generation capability, 56 participants were shown a randomly sampled set of 16 forced-choice questions. Our study consists of four sections corresponding to four different user studies. First section focuses on the comparison with the state-of-the-art approaches. Each question comprises of a side-by-side animation of our method along with one of EMAGE [40], Audio2Photoreal [51], RemoDiffuse [68] or the ground-truth. Second section is a short section where we compare LLM-driven gesture generation with discourse-based synthesis. Third, we perform pair-wise comparisons between results from one of the RAG baselines (derived from ReMoDiffuse [68]) and our method. Lastly, we evaluate different approaches to perform RAG by controlling its extent.
In first three sections, we try to evaluate naturalness and appropriateness. Specifically, we ask two questions (a) “Which of the two gestures look natural?” and (b) “Which of them looks appropriately aligned to what the person is saying?”. In the last section, we add an additional question which focuses on gauging the semantic appropriateness in the retrieval window, with a goal of evaluating the RAG capability. We highlight the identified words from the retrieval algorithms and add an additional question: “Which of the two have better gestures in the highlighted section, especially at the capitalized word in the prompt?”
8 Evaluation Metrics
FID.
Beat Alignment Score.
L1 Divergence.
This metric (also called L1 variance) measures the distance of all frames in a single generated sample from their mean. It is helpful in identifying synthesized gestures that are static and unexpressive.
Diversity.
It computes the average pairwise Euclidean distance between the generated gestures from the test set.
Multi-modality.
This metric requires sampling different gesture motions for a single speech input from the generative model [12]. Then, it computes Euclidean distance between the diverse generated gestures. It probes the diverse sampling capabilities of a generative model.
@@ 8 Evaluation Metrics
Table 5: Overall Multi-modality for All-Speaker Model
9 Additional Experiments
9.1 Per-speaker Quantitative Comparison with FID & Multi-modality
@@
To perform a robust evaluation on speaker generalizability of our framework, we provide per-speaker FID and Multi-modality metrics for the all-speaker model in Tab. 4. We observe that our framework achieves best FID for large number of the speakers, which shows that our method generalizes well to the speaker specific patterns and idiosyncrasies despite taking no seed gestures at input. Overall, CaMN [39], achieves lower FID because it uses seed input from the ground truth data which results in lower scores. Due to the same reason, EMAGE [40] also gets lower FID.
@@
However, CaMN and EMAGE always generate same gestures for a given speech input, so they perform worse in terms of Multi-modality, which makes them less ideal for diverse gesture generation. In contrast, our approach gets the best results showing diverse gesture generation capabilities of our model ( Tab. 5).
9.2 Comparison of RAG with Motion Blending
Linear motion blending can be considered as an alternative for exemplar insertion in the generated motions. Therefore, we explore this alternative by pasting semantic examples onto the retrieval windows and blending motion at the window boundaries. We observe smoothing artifacts with motion blending where motion looks unnatural and gesture beats are smoothened around window boundaries due to interpolation. Compared to this, motion naturalness and speech-to-gesture alignment are preserved by using the proposed retrieval insertion method, which “blends” motion in the diffusion latent space. Video results can be seen in the supplementary video (@12:54).
9.3 Ablation on single speaker training
We observe high FID for the single-speaker setting. Therefore, we analyze this further in Tab. 6 by comparing single-speaker model with Non-RAG version (to check underfitting), and also with an RAG version which uses larger all-speaker database but same model trained on a single speaker (to check if smaller database causes worse performance).
[…]
GT $0.703$ $11.97$ $127$
No RAG $0.911$ $0.727$ $12.78$ $130$
RAG with all-speaker DB (w/ Discourse) $0.872$ $0.727$ $12.53$ $127$
RAG with 1-speaker DB (w/ Discourse) $0.879$ $0.730$ $12.62$ $129$
Table 6: Quantitative Ablation with 1-speaker training.
We observe the same trend as the ablative analysis, that RAG versions of the model perform better. Interestingly, usage of larger DB performs slightly better due to better example matching. However, FID remains higher than CaMN/EMAGE and metrics show little difference between 1-speaker model and “No RAG” model, which shows underfitting due to small data size. Our experiments concur that diffusion models are more data-hungry and train better with larger data. Moreover, we observe that deterministic models (CaMN/EMAGE) that use seed motion, perform better with the smaller data by overfitting and predicting samples closer to GT (Sec. 9.1).
10 Implementation Details & Analysis
@@ 10.1 Input Representations
Representing speech and its transcription is highly important aspect of diffusion-based gesture modelling process. In our experiments, we found that changing the structure of text embeddings affects gesture understanding during the learning process, which consequently is reflected during the synthesis phase as well. To construct our text representation, we build a per-frame embedding with corresponding word embeddings residing on each frame. To extract discourse connectives, we pass text transcriptions of speech samples through discopy [34] and store resulting outputs along with dataset samples. We compute the word embeddings by aggregating sub-word token activations from last 4 layers of BERT model [13].
@@ 10.2 Decoupled gesture encoding
We utilize time-aware VAE architecture by Mughal et al. [47]. This architecture utilizes seperate encoders for frame window chunks of original motion to encode each chunk into an encoding. Then, it jointly decodes all of the chunks together to reconstruct the original motion. We use $N=150$ representing $10$ seconds of motion at $15$ frames per second. Moreover, the frame chunk length of our time-aware VAE is $15$, making each chunk encoding correspond to $1$ second of motion. This results in a chunked gesture encoding of length $10$ for each body part. Finally, we concatenate all 4 body part encodings along the time axis with separators in between them, resulting in $M=40+3=43$.
@@ 10.3 RAG-driven Gesture Diffusion model
In the current setup, the insertion of retrieved gestures happens at $t=T$ where the latents are fully noised. However, our implementation also allows to arbitrarily choose a timestamp $t=K$ for retrieval insertion. Consequently, one can then perform Retrieval Guidance for steps $t<K$.
@@ 10.4 Details on LLM Prompting
“You are an expert in human gestures. You need to identify words that may elicit semantically meaningful gestures(deictic, iconic, metaphoric) and their types: (a) Metaphoric Gesture: Represents abstract ideas or concepts physically, creating a vivid mental image. (b) Iconic Gesture: Mimics the shape or action of the object or concept being described. (c) Deictic Gesture: Points to or indicates a person, object, or location. Format your response as a python list of python tuples of (word, type). For example: [(’hello’, ’beat’), (’world’, ’iconic’)]”.
@@
Identify at most 2 important words which are more likely to elicit semantically meaningful gestures and what are types of those gestures in following text: “SAMPLE TEXT”.
Effect of Word Number in Prompt.
[…]
GT $0.477$ $7.29$ $110$
1-word $0.483$ $0.486$ $9.38$ $115$
2-word $0.487$ $0.514$ $9.94$ $118$
3-word $0.510$ $0.536$ $10.26$ $120$
Table 7: Quantitative Comparison using different quantities of identified words in LLM prompt.
As shown above, the user prompt contains a maximum number of words for which LLM needs to predict gesture types. As this number of words is a hyperparameter, we perform quantitative comparison for different quantities in Tab. 7. Results show slight variation in metrics and even smaller difference in terms of perceptual quality. Therefore, we conclude that retrieval algorithm is flexible enough to be used with any configuration.
@@ 10.5 Baseline Retraining Details
To be consistent with single speaker evaluation on BEAT2 dataset, we utilize released model weights by EMAGE [40]. For other approaches (including ours), we retrain the method on single speaker data belonging to the speaker “Scott”. Since there are no available models for the chosen baselines which have been trained on all speakers in BEAT2 dataset, we train all the methods on complete dataset through their provided codebases. Methods which do not contain speaker specific generalizations like Audio2Photoreal [51], are modified to include a speaker embedding along with text and speech embeddings. Moreover, Audio2Photoreal is adapted to support the skeletal format of BEAT2. Lastly, ReMoDiffuse, originally released for text-to-motion task, is modified for gesture synthesis and their retrieval process is implemented using text feature similarity method.
For the comparison of our approach with training-based RAG (Sec. 4.3), we further modify the ReMoDiffuse architecture and train it using our gesture encodings instead of raw motion. Importantly, we implement retrieval merging strategy of SemanticGesticulator [71] to incorporate the output of our proposed retrieval algorithms into this training-based approach. We perform this experiment by training it on all speakers and utilizing Discourse-based retrieval algorithm.
10.6 Runtime Information:
Retrieval algorithms involve multiple ranking and filtering steps, each contributing to a certain computation time. Specifically, Discourse-based algorithm takes 0.03s to run on 1 data sample. LLM-based algorithm also includes an API call along with ranking steps and therefore, total time taken is increased to 1.52s which includes 0.95s for API call. RAG-driven inference for the diffusion model takes 26.93s on NVIDIA RTX3090 for a batch size of 32.